OpenStudy (sloppycanada):

Just a couple rational functions questions that if people don't mind helping me work through, that'd be great!

OpenStudy (sloppycanada):



OpenStudy (amoodarya):

\[\frac{5x+1}{x}<1\\\frac{5x+1}{x}-1<0\\\frac{5x+1}{x}-\frac{x}{x}<0\\\frac{5x+1-x}{x}<0\] can you go on ?

OpenStudy (amoodarya):

4x+1 sorry

OpenStudy (amoodarya):

can you go on ?

OpenStudy (amoodarya):